# Neural Representations of Dynamic Visual Stimuli

### v1

@@ Abstract
Humans experience the world through constantly changing visual stimuli, where scenes can shift and move, change in appearance, and vary in distance. The dynamic nature of visual perception is a fundamental aspect of our daily lives, yet the large majority of research on object and scene processing, particularly using fMRI, has focused on static stimuli. While studies of static image perception are attractive due to their computational simplicity, they impose a strong non-naturalistic constraint on our investigation of human vision. In contrast, dynamic visual stimuli offer a more ecologically-valid approach but present new challenges due to the interplay between spatial and temporal information, making it difficult to disentangle the representations of stable image features and motion. To overcome this limitation – given dynamic inputs, we explicitly decouple the modeling of static image representations and motion representations in the human brain. Three results demonstrate the feasibility of this approach. First, we show that visual motion information as optical flow can be predicted (or decoded) from brain activity as measured by fMRI. Second, we show that this predicted motion can be used to realistically animate static images using a motion-conditioned video diffusion model (where the motion is driven by fMRI brain activity). Third, we show prediction in the reverse direction: existing video encoders can be fine-tuned to predict fMRI brain activity from video imagery, and can do so more effectively than image encoders. This foundational work offers a novel, extensible framework for interpreting how the human brain processes dynamic visual information.
@@ 1 Introduction
From a car driving down the street to waves crashing against the shore, our daily lives are filled with visual stimuli that are constantly in motion. The dynamic and spatiotemporally rich nature of visual perception is a cornerstone of our daily experience, and the human visual system has evolved to extract meaningful patterns from the dynamics of the visual environment. Yet, dynamics have often been omitted in the study of the neural basis of high-level visual perception, with static images being the primary focus of many studies. Such inputs only capture a snapshot of the visual world, omitting the rich temporal information that is inherent in our everyday experiences. This omission is largely due to the relative computational simplicity of static images, which, while leading to significant advances in our understanding of the neural representations of visual features, has come at the cost of neglecting the dynamic aspects of visual perception. For example, a video of a person walking affords not only information about the visual features of the walker, but also about motion patterns, context, and relationships between the different elements in the scene (e.g., sudden increase in gait implies urgency). This interplay between spatial and temporal information may be impossible to capture from static images. Consequently, the neural responses evoked by dynamic visual stimuli are likely to differ from those evoked by static images, and understanding these differences is a significant challenge in developing a complete theory of visual perception.
In this paper, we propose a novel framework for analyzing dynamic visual stimuli, which explicitly decouples the modeling of static image representations and motion representations in the human brain. By doing so, we can provide a framework for studies that will enable a deeper understanding of how the brain processes dynamic visual stimuli. In particular, while some scene semantics are available from static images, dynamic visual perception is important to understanding ecologically critical domains such as social interaction and action representation. In the former domain, social roles are are often inferred through the relative physical motions of different actors [1] and in the latter domain, different actions often appear the same in static images, for example, opening versus closing a door, or standing up versus sitting down. For both, adding motion removes ambiguity.
@@
Contributions. Our contributions are as follows. First, we demonstrate that visual motion information in the form of optical flow can be predicted (or decoded) from brain activity as measured by fMRI. Second, we show that this predicted motion can be used to realistically animate static images using a motion-conditioned video diffusion model (where the motion is driven by fMRI brain activity). We demonstrate that brain activity can produce more reliable animations than image-conditioned video diffusion models trained on massively-large datasets. Finally, we show prediction in the reverse direction: recent video encoders (such as VideoMAE and VC-1 [2, 3]) can more effectively predict fMRI brain activity from video imagery as compared to image encoders. Overall these results demonstrate the effectiveness of our approach as a novel, extensible framework for interpreting how the human brain processes dynamic visual information.
@@ 2 Related Work
Image decoding from fMRI
Reconstructing visual images from human brain activity, such as that measured by functional magnetic resonance imaging (fMRI), has significantly advanced through the use of deep learning models trained on extensive datasets of natural images [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Recently, the number of approaches that integrate various aspects of visual experiences to reconstruct visual perceptions has increased rapidly [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. This progress is largey due to the availability of large, publicly accessible fMRI datasets with well-annotated stimuli, such as BOLD5000 [29] and the Natural Scenes Dataset (NSD) [30], along with multimodal deep generative models that have been pre-trained on very large datasets, such as Stable Diffusion [31].
Video Decoding from fMRI
While there has been extensive decoding work on image reconstruction from fMRI, there has been much less work on video reconstruction. The most notable example of the latter is perhaps MindVideo [19], which performs pretraining to align temporal semantics from fMRI and simultaneously generate frames across time. Notably, our approach explicitly disentangles the representations of motion information (represented as optical flow) from static images in fMRI signals. This allows us to differentiate our present work from previous work by enabling interpretability. Specifically, we isolate and focus on decoding motion (as optical flow) from fMRI signals, as opposed to more difficult to interpret mixture of static image features and motion. This disentanglement allows us to learn an explicit representation of motion which can be directly compared to groundtruth and be more directly related to spatially localized neural mechanisms, as opposed to an entangled mixture of static image and dynamic features.
@@ 3.1 Motion Prediction
Our primary objective is to extract the information relevant to visual motion in order to predict the optical flow in videos observed by subjects. To enhance interpretability, we’ve architecturally separated static image features from dynamic features. By conditioning the model on the initial frame alongside neural data, we enable our model to focus on extracting dynamic visual information. This approach improves interpretability compared to models like MindVideo, which simultaneously predict both dynamic and static features. In contrast, our approach disentangles the specific contributions of dynamic components.
At a given fMRI scan at index we have fMRI voxels and corresponding video frames . We obtain optical flow fields from RAFT [32], which consist of optical flow vectors in . Prior work [33] has found that reformulating optical flow prediction as classification improves performance. We take a similar approach and quantize the optical flow vectors into a codebook of clusters using k-means. We obtain image features with pretrained vision model . More concretely, we learn a function that classifies the fMRI voxels and image features of the corresponding initial frame to the quantized future predicted optical flows fields:
@@
Visualizing the Predicted Motion. We can visualize the predicted motion by passing in the predicted optical flow field and the initial frame to a pretrained motion-conditioned video diffusion model, DragNUWA [35]. This allows us to generate the dynamic vision stimuli (observed by the subject) or (driving the neural response).
3.2 Visual Stimuli-To-Brain Encoding.
@@
Semantically-supervised: CLIP (ViT-B-32) and CLIP ConvNeXt are trained to learn align semantics and images. CLIP uses a visual transformer architecture and CLIP ConvNeXt uses a convolutional backbone, sharing inductive biases with ResNet-50 but trained with a different objective. CLIP AF and CLIP ConvNeXt AF consist of CLIP and CLIP ConvNeXt features averaged over frames spanning the video [39, 40].
@@ 4.1 Dataset
We utilize the Dynamic Natural Vision dataset [45], a public fMRI dataset composed of brain responses from 3 participants viewing diverse naturalistic video clips. The dataset contains 11.5 hours of fMRI data, with the 3 subjects viewing the same videos. The fMRI data was recorded using a 3T MRI system with a 2 second TR and the videos were presented at 30 FPS. 18 8-minute videos were seen twice and 5 8-minute videos were seen 10 times, including the train and test sets which have no overlapping videos. We use the same train/test split described in [45]. The fMRI data is preprocessed and aligned with presented stimuli using the pipeline from the Human Connectome Project. We further normalized the fMRI activations so that each voxel’s response was zero-mean and unit variance on a session basis. Then we average the response across repeated viewings of the same visual stimuli.
@@
fMRI captures the blood oxygen level dependent (BOLD) signal, which peaks approximately 4 seconds after stimulus presentation and slowly decays after, due to nature of the hemodynamic response. During preprocessing, the fMRI data is shfited to temporally align the peak of the response to the onset of its corresponding visual stimuli presentation. However, due to the long time scale of the fMRI response, fMRI scans 4 seconds or 2 TRs prior and after the current TR may have information shared with the current TR. In order to leverage all relevant neural responses, when predicting the motion in the video presented at TR=, we concatenate fMRI data from TRs . We generate the video frames conditioned on the predicted motion using DragNUWA [35].
@@ 4.2 Motion Decoding
Quantitative Flow Evaluation.
First we report evaluations of the decoded optical flow vectors on both direction and end point error on the test set. We quantitatively compare the models trained on the 3 subjects against a baseline of the motion decoder that does not use neural data and only image features from DINOv2 in Figure 3. Since this baseline model is not conditioned on neural data, the model learns to regress plausible flows which do not necessarily align with the true flow. Indeed, we observe that the models trained on neural data perform significantly better on both end point error ([] vs , for the three models trained on neural data) and cosine similarity ([] vs , . In addition, across the testing dataset, we find that models trained on neural data predict directions that are positively aligned with the true flow directions. In contrast, the model trained without neural data predicts flow directions that are actually anti-aligned with the true flow directions.
Reasoning With and Without fMRI.
@@
Here we present examples and explain why the models trained on neural data outperform the model trained without neural data (Figure 4).
@@
First, there exists ambiguity of action given static images - a frame may contain multiple possible actions. For instance, in Figure 4a we show an eagle flying. Given the first frame, the eagle could be plausibly either flapping its wings or gliding. The model trained on neural data is able to identify that the eagle is gliding. However, the model trained without neural data incorrectly predicts that the eagle is flapping its wings. Second, even if the action is not ambiguous, there exists camera motion in the videos which can change the direction of the flow fields. For example, in Figure 4b we show a space shuttle flying into the sky. In this case, with no camera motion the direction of motion is clearly upwards and to the right, which the model trained without neural data correctly predicts. However, because the video is centered on the space shuttle with camera motion, the exhaust from the space shuttle appears to be going downwards and to the left. This causes the optical flow vectors to point downwards and to the left. Again, the model trained with neural data is able to correctly predict the motion of the object. Third, there exists ambiguity of direction of motion in static images of stationary objects. For instance, if the first frame contains just a stationary portion of the Eiffel Tower, what movement should be predicted? One could choose to predict no motion, or arbitrarily predict a direction of camera motion. Indeed, we see in Figure 4c that the trained model without neural data chooses an arbitrary camera motion pointing leftwards. However, the model trained with neural data correctly predicts the true direction which is camera panning downwards.
@@
The results derived from these scenarios indicate that the neural data contains information about the correct movement which cannot be derived from just image features.
@@
We show additional motion decoding results along with the video reconstruction from DragNUWA (Figure 5). We find that the motion we predict from the fMRI brain activity is not only aligned with the ground truth motion, but also generates realistic videos. We further demonstrate that we are able to decode motion that is consistent and that suddenly changes direction, i.e. the airplane example in Figure 5.

### v2

# Reanimating Images using Neural Representations of Dynamic Stimuli

@@ Abstract
While computer vision models have made incredible strides in static image recognition, they still do not match human performance in tasks that require the understanding of complex, dynamic motion. This is notably true for real-world scenarios where embodied agents face complex and motion-rich environments. Our approach leverages state-of-the-art video diffusion models to decouple static image representation from motion generation, enabling us to utilize fMRI brain activity for a deeper understanding of human responses to dynamic visual stimuli. Conversely, we also demonstrate that information about the brain’s representation of motion can enhance the prediction of optical flow in artificial systems. Our novel approach leads to four main findings: (1) Visual motion, represented as fine-grained, object-level resolution optical flow, can be decoded from brain activity generated by participants viewing video stimuli; (2) Video encoders outperform image-based models in predicting video-driven brain activity; (3) Brain-decoded motion signals enable realistic video reanimation based only on the initial frame of the video; and (4) We extend prior work to achieve full video decoding from video-driven brain activity. This framework advances our understanding of how the brain represents spatial and temporal information in dynamic visual scenes. Our findings demonstrate the potential of combining brain imaging with video diffusion models for developing more robust and biologically-inspired computer vision systems. We show additional decoding and encoding examples on this site: https://sites.google.com/view/neural-dynamics/home.
@@ 1 Introduction
Human vision excels at understanding complex, dynamic scenes, while computer vision offers powerful tools for data processing and analysis. By combining these complementary strengths, we can develop more robust and adaptable vision systems [5]. This work explores how insights from human brain activity, specifically in response to dynamic visual stimuli, can enhance the ability of computer vision systems to understand complex motion. Critically, as inherently embodied systems, our brains have evolved to process and represent the daily experience of naturally motion-rich environments in ways that are both semantically grounded and behaviorally-relevant. For example, a video of a person walking affords information not only about the visual features of the walker, but also about motion patterns, context, intent, and relationships between the different elements in the scene (e.g., sudden increase in gait implies urgency). This interplay between spatial and temporal information is impossible to capture from static images; accordingly, humans excel at inferring such high-level information from scene dynamics [6].
While recent work has made significant strides in understanding the neural representation of visual information using computer vision models [7, 8] there remains a notable gap in our understanding of how the brain processes dynamic visual stimuli. Although some studies have begun to investigate the neural representation of dynamic visual stimuli [3, 9, 10], there is still a vast opportunity to leverage modern computer vision techniques, particularly video-based models. Second there has been little work demonstrating how insights from human vision can be leveraged to build better computer vision systems. Here we address both challenges, demonstrating that state-of-the-art video foundation models can be used to robustly predict human brain responses to dynamic visual inputs and that brain-decoded motion signals can help to reanimate videos based only on the first frame of the video. Critically, our framework outperforms prior work for both problems, achieving lower end point error (EPE) for motion decoding based on brain responses as compared to [3] and better motion prediction based on initial static frames with brain data than without brain data.
In this paper, we present a framework for analyzing dynamic visual stimuli that explicitly separates the modeling of static and motion-related image representations as they are processed in the human brain. By decoupling these elements, our framework offers a targeted approach to understanding the distinct neural pathways used to process and represent dynamic visual information in real-world scenes.
While an important question in its own right, better specified models of dynamic visual processing are also critical for developing effective models of ecologically critical domains such as social interaction, action representation, and dynamic causal structure – none of which are easily extracted from static images. For example, social roles are are often inferred through the relative physical motions of different actors [6]. Similarly, many different actions appear the same in static images, for instance, opening versus closing a door, or standing up versus sitting down. Finally, learning about functional world knowledge (e.g., tool use or context-appropriate physical interactions) is primarily acquired through observation and imitation of dynamic events [11]. For these and many other domains, adding motion removes ambiguity.
@@
Immediate Contributions. In this context, our project makes several immediate contributions. First, we demonstrate that visual motion information in the form of optical flow can be predicted (or decoded) from brain activity as measured by fMRI. Second, we show that this predicted motion can be used to realistically animate static images using a motion-conditioned video diffusion model in which the motion is driven by brain activity. Moreover, we demonstrate that brain activity can produce more reliable animations than image-conditioned video diffusion models trained on massively-large datasets. Third, we apply our method to existing methods that decode the viewed video by using their initial frame to do full video decoding from video-driven brain activity and find that we more accurately predict the viewed motion than existing methods. Finally, we show prediction in the reverse direction: recent video encoders (such as VideoMAE and VC-1 [1, 12]) can more effectively predict fMRI brain activity from video imagery as compared to image encoders. Overall these results demonstrate the effectiveness of our approach as a novel, extensible framework for interpreting how the human brain processes dynamic visual information.
Larger Objectives. Our project also has several larger objectives. First, developing high-performing models for motion decoding based on neural data (Figure 1d) is an important precursor to building effective encoding models [13]. Second, the finding that VideoMAE is able to accurately predict brain responses (Figure 1a) is an important step towards building more general and informative video encoding models for brain prediction [8]). Encoding models are also important components in comparisons between model-based representations and empirical neural data, which can reveal key insights into which visual features the brain prioritizes and how they align with computational processing in interpretable artificial neural networks [14]. Third, we apply our method to existing methods which decode the viewed video to do full video decoding from video-driven brain activity [3] and find that we more accurately predict the viewed motion than existing methods. Fourth, although comparison points are limited due to the underexploration of this research area, our finding that neural data improves video prediction performance better than other extant models is a simple demonstration of the point that latent information encoded in brain data may be useful for improving the performance of computer vision systems [15]. This approach may prove to be especially powerful for domains that are inherently anchored in human behavior, for example, dynamic social interactions.
@@ 2 Related Work
Static Visual Stimuli Decoding from fMRI. Reconstructing visual images from human brain activity, such as that measured by functional magnetic resonance imaging (fMRI), has significantly advanced through the use of deep learning models trained on extensive datasets of natural images [7, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Recent methods leverage deep generative models to analyze and decode brain representations [32, 33, 3, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. This progress is largely due to the availability of large, publicly accessible fMRI datasets with well-annotated stimuli, such as BOLD5000 [59] and the Natural Scenes Dataset (NSD) [60], along with multimodal deep generative models that have been pre-trained on very large datasets, such as Stable Diffusion [61].
Dynamic Visual Stimuli Decoding from fMRI. While there has been extensive decoding work on image reconstruction from fMRI, there has been much less work on video reconstruction. Within the work on decoding motion from fMRI, much of prior work focuses on classifying global motion in artificial stimuli like random dots moving in a coherent direction [62, 63]. Recent work has also explored disentangling static and motion features using semantically meaningful dot movements [64]. We significantly advance this by decoding motion in complex, naturalistic visual stimuli at a more fine-grained, object-level resolution. This allows us to study how the brain represents motion in more ecologically valid scenarios.
Our fMRI-based approach offers a whole-brain perspective compared to single-neuron recordings, which are often limited in spatial coverage [65, 66, 67, 62, 63, 68]. Despite the challenges of fMRI’s relatively slow temporal resolution, our results demonstrate successful decoding of motion information from complex, naturalistic stimuli at the object level. This suggests that fMRI, despite its limitations, can still provide valuable insights into the neural representation of dynamic information.
Concurrent to motion decoding research, there is a parallel branch of research focusing on decoding the raw video from fMRI activity [71, 72, 3, 73, 74, 75]. Methods like MindVideo [3] train to align temporal semantics from fMRI and generate frames across time. Notably, our approach explicitly disentangles the representations of motion information (represented as optical flow) from static images in fMRI signals. This allows us to differentiate our present work from previous work by enabling interpretability. Specifically, we isolate and focus on decoding motion (as optical flow) from fMRI signals, as opposed to more difficult to interpret mixture of static image features and motion. This disentanglement allows us to learn an explicit representation of motion which can be directly compared to groundtruth and be more directly related to spatially localized neural mechanisms, as opposed to an entangled mixture of static image and dynamic features.
@@ 3.1 Motion Prediction
Our primary objective is to extract the information relevant to visual motion in order to predict the optical flow in videos observed by participants. To enhance interpretability, we’ve architecturally separated static image features from dynamic features. By conditioning the model on the initial frame alongside neural data, we enable our model to focus on extracting dynamic visual information. This approach improves interpretability compared to models like MindVideo, which simultaneously predict both dynamic and static features. In contrast, our approach disentangles the specific contributions of dynamic components.
At a given fMRI scan at index we have fMRI voxels and corresponding video frames . We obtain optical flow fields from RAFT [76], which consist of optical flow vectors in . Prior work [77] has found that reformulating optical flow prediction as classification improves performance. We take a similar approach and quantize the optical flow vectors into a codebook of clusters using k-means. We obtain image features with pretrained vision model . Concretely, we learn a function that classifies the fMRI voxels and image features of the corresponding initial frame to the quantized future predicted optical flows fields: .
@@
Visualizing the Predicted Motion. We can visualize the predicted motion by passing in the predicted optical flow field and the initial frame to a pretrained motion-conditioned video diffusion model, DragNUWA [4]. This allows us to generate the dynamic vision stimuli (observed by the participants) or (driving the neural response).
3.2 Visual Stimuli-To-Brain Encoding
@@
Semantically-supervised: CLIP ViT-B/32 and CLIP ConvNeXt are trained to learn align semantics and images. CLIP ViT-B/32 uses a visual transformer architecture and CLIP ConvNeXt uses a convolutional backbone, sharing inductive biases with ResNet-50 but trained with a different objective. CLIP ViT-B/32 AF and CLIP ConvNeXt AF consist of features Averaged over Frames (AF) spanning the video [81, 82].
@@ 4.1 Dataset
We utilize the Dynamic Natural Vision dataset [87], a public, IRB approved fMRI dataset composed of brain responses from 3 participants viewing diverse naturalistic video clips. The dataset contains 11.5 hours of fMRI data, with the 3 participants viewing the same videos. The fMRI data was recorded using a 3T MRI system with a 2 second TR and the videos were presented at 30 FPS. 18 8-minute videos were seen twice and 5 8-minute videos were seen 10 times, including the train and test sets which have no overlapping videos. We use the same train/test split described in [87]. The fMRI data is preprocessed and aligned with presented stimuli using the pipeline from the Human Connectome Project. We further normalized the fMRI activations so that each voxel’s response was zero-mean and unit variance on a session basis. Then we average the response across repeated viewings of the same visual stimuli.
@@
fMRI captures the blood oxygen level dependent (BOLD) signal, which peaks approximately 4 seconds after stimulus presentation and slowly decays after, due to nature of the hemodynamic response. During preprocessing, the fMRI data is shifted to temporally align the peak of the response to the onset of its corresponding visual stimuli presentation. However, due to the long time scale of the fMRI response, fMRI scans 4 seconds or 2 TRs prior and after the current TR may have information shared with the current TR. In order to leverage all relevant neural responses, when predicting the motion in the video presented at TR=, we concatenate fMRI data from TRs . We generate the video frames conditioned on the predicted motion using DragNUWA [4].
@@ 4.2 Motion Decoding
Quantitative Flow Evaluation. First we report evaluations of the decoded optical flow vectors on the end point error on the test set. We quantitatively compare the models trained on the 3 participants against baselines of: MindVideo (Best) [3], our method leveraging the initial frame generated from MindVideo (Ours + MindVideo (Best)), and Stable Video Diffusion [2] that does not use neural data and only the initial frame in Figure 3. We extract the optical flows from the videos generated by MindVideo and Stable Video Diffusion using the same RAFT model we use to extract the ground truth flow. Since MindVideo and Stable Video Diffusion are generative models that can produce multiple plausible outputs, we generate 100 samples per video for MindVideo and 10 samples for Stable Video Diffusion (because the latter is able to condition on the true video’s first frame) and select the sample with the lowest end point error for fair comparison. We evaluate our method using the true initial frame from the video (Ours) as well as the first frame generated by MindVideo (Ours + MindVideo(Best)). Since the Stable Video Diffusion baseline (SVD (Best)) is not conditioned on neural data, it learns to predict plausible flows which do not necessarily align with the true flow. Indeed, we observe that our model trained on individual participant neural data performs significantly better on end point error as compared to SVD (Best) (end point error for individual participants – S1, S2, S3 – respectively, S1: 0.543, S2: 0.572, S3: 0.634 vs. 1.192, all using paired -tests with Holm-Bonferroni correction).
@@
Reasoning With and Without fMRI. We present examples and explain why our models trained on neural data outperform the model without brain conditioning, Stable Video Diffusion (Figure 4).
First, there exists ambiguity of action given static images - a frame may contain multiple possible actions. For instance, in Figure 4a we show an eagle flying. Given the first frame, the eagle could be plausibly either flapping its wings or gliding. The model trained on neural data is able to identify that the eagle is gliding. However, the model without brain conditionining incorrectly predicts that the eagle is flapping its wings.
@@
Second, even if the action is not ambiguous, there exists camera motion in the videos which can change the direction of the flow fields. For example, in Figure 4b we show a space shuttle flying into the sky. In this case, with no camera motion the direction of motion is clearly upwards and to the right, which the model without brain conditioning correctly predicts. However, because the video is centered on the space shuttle with camera motion, the exhaust from the space shuttle appears to be going downwards and to the left. This causes the optical flow vectors to point downwards and to the left. Again, the model trained with neural data is able to correctly predict the motion of the object.
Third, there exists ambiguity of direction of motion in static images of stationary objects. For instance, if the first frame contains just a stationary portion of the Eiffel Tower, what movement should be predicted? One could choose to predict no motion, or arbitrarily predict a direction of camera motion. Indeed, we see in Figure 4c that the model without brain conditioning chooses an arbitrary camera motion pointing leftwards. However, the model trained with neural data correctly predicts the true direction which is camera panning downwards. We provide additional controls and quantitative results in the supplemental.
@@
Initial Frame Reanimation. The results derived from these scenarios indicate that the neural data contains information about the correct movement which cannot be derived from just image features.
@@
We show motion decoding results along with the video reconstruction from DragNUWA (Figure 5). We find that the motion we predict from the fMRI brain activity is not only aligned with the ground truth motion, but also generates realistic videos. We further demonstrate that we are able to decode motion that is consistent and that suddenly changes direction, i.e. the airplane example in Figure 5 .
Because we disentangle the modeling of static image and motion representations, we can extend our method to other works that decode image frames from fMRI. We demonstrate examples of extending our method to initial frames decoded from MindVideo in Figure 6 and find we decode motion more accurately (Figure 3). We provide additional video examples in the supplemental.